# Manifolds with kinks and the asymptotic behavior of the graph Laplacian operator with Gaussian kernel

### v2

@@ 1. Introduction
The connection between the graph Laplacian and the Laplace–Beltrami operator is a central theme in geometric data analysis and has attracted substantial interest from both applied and theoretical communities. On the one hand, from a theoretical perspective, this relationship offers a discretization framework for studying differential operators on manifolds via point clouds or sampled data [LSS17, HSAY20, YJH23, JYHL24]. On the other hand, from an applied viewpoint, this connection underpins a wide array of algorithms in machine learning and signal processing, including manifold learning techniques such as Laplacian eigenmaps [BN03, BN05] or Diffusion Maps [CL06, SW12]. Understanding the convergence of the graph Laplacian to the Laplace–Beltrami operator — notably the role of scaling, sampling density, and boundary behavior — remains crucial in ensuring that discrete approximations faithfully capture the geometry and analysis of the underlying continuous space. This paper aims at tackling this question on a class of smooth Riemannian manifolds having possibly singular boundary.
@@
After the work of various authors, e.g. [HAvL05, BN05], it is by now known that when $M$ is a smooth $d$-dimensional submanifold of an Euclidean space $\mathbb{R}^{D}$, for any sufficiently regular $f$ and $p$, the extrinsic Gaussian operator satisfies
@@
with $\Delta_{M}$ the classical Laplace–Beltrami operator of $M$. In [BQWZ12] (see also [AA25]), this pointwise result for the extrinsic Gaussian operator was extended to three types of singular subsets of $\mathbb{R}^{D}$, namely submanifolds with boundaries, intersections of them, and submanifolds with edge-type singularities. For instance, if $x$ belongs to the non-empty boundary of some smooth submanifold $M\subset\mathbb{R}^{D}$, then
@@
Our main results extend both (1) and (2) in the setting of smooth Riemannian manifolds with kinks. These spaces, which we introduce in Section 3, might be understood as manifolds with possibly non-empty and irregular boundary. To distinguish these possibly non-smooth boundaries with classical smooth ones, we use the terminology border : a smooth manifold with kinks $M$ thus divides into an interior part $intM$ and a border part $\partial M$. Manifolds with kinks notably include manifolds without boundary, manifolds with smooth boundary, and manifolds with corners as considered e.g. in [Cer61, Dou61, DH73, Mic80, Mel96, Joy12]. Other relevant spaces like cones, pyramids, even cusps, fall into the framework of manifolds with kinks while they are not manifolds with corners.
@@
For the purpose of our analysis, we need minimal regularity on the boundary of the manifolds with kinks that we consider. Thus we work under Lipschitz and Continuously Directionally Differentiable (LCDD for short) regularity, noting that this condition is satisfied by a large class of manifolds with kinks, e.g. manifolds with boundaries and corners, and pyramids. Our approach also allows for cusps, which are not LCDD. We refer to Section 3 for the precise definitions of LCDD and cusps in this context.
@@
$L_{t}f(x)$ $=-\frac{c_{d}}{\sqrt{t}}\left(p(x)\,\partial_{v_{g}(x)}f(x)+o(1)\right)-c_{d+1}\bigg{(}p(x)A_{g}f(x)+[p,f]_{g}(x)\bigg{)}+O(\sqrt{t})+O(t^{-d}e^{-t^{2\eta-1}})$
as $t\downarrow 0$, for any $f\in\mathcal{C}^{3}(M)\cap L^{1}(M,p\,\mathrm{vol}_{g})$, where $\partial_{v_{g}(x)}f(x):=\mathop{}\!\mathrm{d}_{x}f\left(v_{g}(x)\right)$ and
@@
$L_{t}f(x)$ $=\left[-\frac{c_{d}}{\sqrt{t}}\left(\tilde{p}(0_{d})\,\partial_{M}\tilde{f}(0_{d})+o(1)\right)-c_{d+1}\bigg{(}\tilde{p}(0_{d})A_{M}\tilde{f}(0_{d})+[\tilde{p},\tilde{f}]_{M}(0_{d})\bigg{)}\right](1+O(1))$
@@
It is worth pointing out that in the previous context, the inward tangent sector $I_{x}M$ coincides with the Bouligand tangent cone of $\Omega$ at $0_{d}$, so that Theorem 1.2 matches up with Theorem 1.1. Let us also stress out that, while bounded geometry of the boundary was assumed in [HAL07], we do not make any assumption of that kind here.
The recent preprint [ART25] provides a result that goes in the direction of Theorem 1.2, but with less details on the differential operators involved there. Our approach allows for a neat identification of the latter in several cases, including the following ones.
@@
Theorems 1.1 and 1.2 open the door to a natural Neumann boundary condition on (sub)manifolds with kinks, that would be $\partial_{v_{g}(x)}f(x)=0$ for any $x\in\partial M$. A natural problem to tackle after that is the spectral convergence of the Graph Laplacian with Gaussian kernel generated from finitely many samples $\{X_{1},\ldots,X_{n}\}$ to the suitable Neumann Laplacian as sample size goes to infinity. We refer to [BN06, SW17, GTGHS20] for related results on manifolds without and with boundary, see also [DT24] for a preliminary investigation with the symmetrized AMV kernel.
@@
We obtain Theorem 1.3 in Section 6 through concentration estimates. Such estimates — studied in e.g. [HAvL05, BQWZ12, AA25] — quantify the error in approximating the Gaussian operator by the graph Laplace operator. Our results generalize this line of work in two key directions :
@@
    We work in the setting of Riemannian manifolds with kinks, whose singularities can be significantly more severe than those of manifolds with boundary considered in [BQWZ12].
@@ Acknowledgments.
The authors are funded by the Research Foundation – Flanders (FWO) via the Odysseus II programme no. G0DBZ23N. They both thank Dominic Joyce for helpful answers concerning manifolds with corners, and Iosif Pinelis who suggested the proof of (6).
@@ 2. Euclidean models
For a subset $A$ of the Euclidean space $\mathbb{R}^{d}$, we let $\bar{A}$ be the closure of $A$ and $int(A)$ be its interior. We recall that a rigid motion of $\mathbb{R}^{d}$ is a linear map $y\mapsto Ay+b$ where $A\in O(d)$ and $b\in\mathbb{R}^{d}$. The epigraph and strict epigraph of $\gamma:\mathbb{R}^{d-1}\to\mathbb{R}$ are respectively defined as
@@
whenever the limit exists. Lastly, the characteristic function of a subset $A$ of a set $X$ will always be denoted by $1_{A}$.
@@ 2.1. Boundary points
For an open set $\Omega\subset\mathbb{R}^{d}$ with a non-empty boundary $\partial\Omega$, we will call boundary points of $\Omega$ the elements in $\partial\Omega$. Inspired by [Eva22, p.626], we define the regularity of boundary points as follows.
@@
    We say that $x$ is a Continuously Directionally Differentiable (CDD for brevity) boundary point if it is a DD point with a function $\gamma$ whose directional derivative $\gamma^{\prime}(x^{\prime};\cdot)$ is continuous on $\mathbb{R}^{d-1}$.
As customary, we may write smooth instead of $\mathcal{C}^{\infty}$, and Lipschitz instead of $\mathcal{C}^{0,1}$. We will essentially work with boundary points that are $\mathcal{C}^{0}$, $\mathcal{C}^{1}$, Lipschitz, or CDD. The combination of Lipschitz and CDD shall be abbreviated to LCDD.
@@ 2.2. Diffeomorphisms and inward sectors
For $k$ a positive integer, we shall use the following notion of $\mathcal{C}^{k}$ diffeomorphism between possibly non-open subsets of $\mathbb{R}^{d}$, see e.g. [Mel96, Section 1.5] for details.
@@ 2.3. Boundary cones of open Euclidean subsets
In this section, we provide classical notions from convex analysis, see e.g. [RW09] for more details. We recall that $C\subset\mathbb{R}^{d}$ is called a cone if it is invariant under multiplication by a positive number, i.e. if $\lambda v\in C$ for any $v\in C$ and $\lambda>0.$
@@
The Bouligand tangent cone is also known as tangent cone or contingent cone in the literature. It is a closed set. One can easily check that our definition is equivalent to the one given in [RW09, Def. 6.1]. Note also that
@@ 2.4. Non-fluctuating boundary and local blow-ups
(a) Fluctuating $\partial\Omega_{1}$
(b) Non-fluctuating $\partial\Omega_{4}$
@@ 3.1. Topological manifolds with kinks
Recall that a topological space is called paracompact if any open cover has a locally finite open refinement, and that a Hausdorff topological space is paracompact if and and only if it admits partitions of unity subordinate to any open cover. For two topological spaces $X,Y$ and a subset $A\subset X$, a map $\varphi:A\to Y$ is a topological embedding if it is an homeomorphism onto its image.
@@ 3.2. Differentiable manifolds with kinks
    We say that $x\in\partial M$ is a $\mathcal{C}^{1}$-boundary point of $M$ if there exists a local border chart $(U,\phi)$ around $x$ such that $\phi(x)$ is a $\mathcal{C}^{1}$ boundary point of $int(\phi(U))$. Otherwise we say that $x$ is an essential kink.
@@
    We say that an essential kink $x\in\partial M$ is an essential corner of depth $k\in\{2,\ldots,d\}$ if there exists a border chart $(U,\phi)$ centered at $x$ such that, up to a rigid motion, the image $\phi(U)$ locally writes as the epigraph of the function $\gamma:(x_{1}\dots x_{d-1})\mapsto\sum_{i=1}^{k-1}|x_{i}|$.
@@
    The pyramid $M:=\{(x,y,z)\in\mathbb{R}^{3}:z^{2}\leq x^{2}+y^{2},z>0\}$ is a manifold with kinks that is not a manifold with corner, because the point $0_{3}\in\partial M$ is an essential kink which is not a corner.
@@
(a) Tip of a pyramid
(b) Cusp
@@ 3.3.1. Tangent space
From now on, we work only with smooth manifolds with kinks, i.e. those manifolds with kinks for which the transition maps are all $\mathcal{C}^{\infty}$. This is to make sure that the tangent space defined below as the space of smooth derivations (i.e. linear forms on the space of smooth functions) is finite dimensional. Indeed, for any $\mathcal{C}^{k}$ manifold with $k<+\infty$, the space of $\mathcal{C}^{k}$ derivations is infinite dimensional : see e.g. [AMR12, Theorem 4.2.41].
In order to introduce smooth derivations on manifolds with kinks, we must first define what a smooth function is. This is what we do in the next definition, largely inspired by the context of manifolds with corners (see e.g. [Joy12, Definition 2.2]).
@@
If $x\in\partial M$ and $f\circ\phi^{-1}:\phi(\bar{U})\cap\mathbb{B}_{\rho}^{d}\to\mathbb{R}$ admits Taylor polynomials of arbitrary high order, then the existence of a smooth extension of $f\circ\phi^{-1}$ to $\mathbb{R}^{d}$ is ensured by Whitney’s extension theorem [Whi34, Theorem I], see also [Ste70, VI.2.]. Note that in the setting of manifolds with corners, the latter upgrades into the well-posedness of Seeley’s linear extension operator [See64], see [Mel96, Section 1.4].
@@
We can now define the differential of a smooth function between smooth manifolds with kinks. There is no change with the case of classical manifolds, see e.g. [Lee12, p.55].
@@
When $x$ is an interior point, a $\mathcal{C}^{1}$ boundary point or an essential corner, this is already known, see [Joy12]. Let us then assume that $x$ is an essential kink which is not an essential corner. Consider a local chart $(U,\phi)$ centered at $x$. Like for manifolds without boundary, we obtain from the chain rule applied to the identities $\phi\circ\phi^{-1}=\mathrm{id}_{\phi(U)}$ and $\phi^{-1}\circ\phi=\mathrm{id}_{U}$ that the differential $\mathop{}\!\mathrm{d}_{x}\phi:T_{x}M\to T_{0_{d}}\phi(\bar{U})$ is a linear isomorphism with inverse $\mathop{}\!\mathrm{d}_{0_{d}}\phi^{-1}$. But the space of derivations at the $\mathcal{C}^{0}$ boundary point $0_{d}$ of $\phi(U)$ coincides with $\mathbb{R}^{n}$, since a basis of this space is given by the classical partial differential operators $\partial\cdot/\partial x_{1},\ldots,\,\partial\cdot/\partial x_{n}$ defined on the open set $\phi(U)$ and naturally extended to any open set $V\subset\mathbb{R}^{d}$ containing $\overline{\phi(U)}$. ∎

### v3

@@ 1. Introduction
The connection between the graph Laplacian and the Laplace–Beltrami operator is a central theme in geometric data analysis and has attracted substantial interest from both applied and theoretical communities. On the one hand, from a theoretical perspective, this relationship offers a discretization framework for studying differential operators on manifolds via point clouds or sampled data [LiShiSun, HarlimSanzYang, YanJiangHarlim, JiaoYanHarlimLu]. On the other hand, from an applied viewpoint, this connection underpins a wide array of algorithms in machine learning and signal processing, including manifold learning techniques such as Laplacian eigenmaps [BelkinNiyogi2003, BelkinNiyogi2005] or Diffusion Maps [CoifmanLafon2006, SingerWu]. Understanding the convergence of the graph Laplacian to the Laplace–Beltrami operator — notably the role of scaling, sampling density, and boundary behavior — remains crucial in ensuring that discrete approximations faithfully capture the geometry and analysis of the underlying continuous space. This paper aims at tackling this question on a class of smooth Riemannian manifolds having possibly singular boundary.
@@
After the work of various authors, e.g. [Hein2005, BelkinNiyogi2005], it is by now known that when $M$ is a smooth $d$-dimensional submanifold of an Euclidean space $\mathbb{R}^{D}$, for any sufficiently regular $f$ and $p$, the extrinsic Gaussian operator satisfies
@@
with $\Delta_{M}$ the classical Laplace–Beltrami operator of $M$. In [belkin+] (see also [AnderssonAvelin]), this pointwise result for the extrinsic Gaussian operator was extended to three types of singular subsets of $\mathbb{R}^{D}$, namely submanifolds with boundaries, intersections of them, and submanifolds with edge-type singularities. For instance, if $x$ belongs to the non-empty boundary of some smooth submanifold $M\subset\mathbb{R}^{D}$, then
@@
Our main results extend both (1) and (2) in the setting of smooth Riemannian manifolds with kinks. These spaces, which we introduce in Section 3, might be understood as manifolds with possibly non-empty and irregular boundary. To distinguish these possibly non-smooth boundaries with classical smooth ones, we use the terminology border : a smooth manifold with kinks $M$ thus divides into an interior part $intM$ and a border part $\partial M$. Manifolds with kinks notably include manifolds without boundary, manifolds with smooth boundary, and manifolds with corners as considered e.g. in [Cerf, Douady, DouadyHerault, Michor, Melrose, Joyce]. Other relevant spaces like cones, pyramids, even cusps, fall into the framework of manifolds with kinks while they are not manifolds with corners.
@@
For the purpose of our analysis, we need minimal regularity on the boundary of the manifolds with kinks that we consider. Thus we work under Lipschitz and Continuously Directionally Differentiable (LCDD for short) regularity, noting that this condition is satisfied by a large class of manifolds with kinks, e.g. manifolds with boundaries and corners, and pyramids. We get refined results under Lipschitz and Twice Continuously Directionally Differentiable (LTCDD for short). Our approach also allows for cusps, which are not LCDD. We refer to Section 2 for the precise definitions of L(T)CDD boundary points and cusps for local Euclidean models, and to 3 for these definitions in the manifold context.
@@
$L_{t}f(x)$ $=-\frac{c_{d}}{\sqrt{t}}\bigg(p(x)\,\partial_{v_{g}(x)}f(x)+\mathrm{Err}(t)\bigg)-c_{d+1}\bigg(p(x)A_{g}f(x)+[p,f]_{g}(x)\bigg)+O(\sqrt{t})+O(t^{-d}e^{-t^{2\eta-1}})$
as $t\downarrow 0$, for any $f\in\mathcal{C}^{3}(M)\cap L^{1}(M,p\,\mathrm{vol}_{g})$, where $\mathrm{Err}(t)=o(1)$ as $t\downarrow 0$, and
$\partial_{v_{g}(x)}f(x)$ $:=\mathop{}\!\mathrm{d}_{x}f\left(v_{g}(x)\right),$
@@
Moreover :
  1. (1)
    If $x$ is an interior point, then $\mathrm{Err}(t)=0$ for any small enough $t>0$.
  2. (2)
    If $x$ is LTCDD, then for any small enough $t>0$,
    (3) $\mathrm{Err}(t)\leq\,4\Gamma(d+2)\,p(x)\,(\|d_{x}f\|_{g_{x}}+\|d_{x}^{(2)}f\|_{g_{x}})\,SC(x).$
    where $SC(x)$ is the total slicewise curvature of $\partial\Omega$ at $x$, defined in (38). In particular, if the total slicewise curvature vector is zero, then $\mathrm{Err}(t)=0$ for any small enough $t>0$.
@@
$L_{t}f(x)$ $=\left[-\frac{c_{d}}{\sqrt{t}}\left(\tilde{p}(0_{d})\,\partial_{M}\tilde{f}(0_{d})+\mathrm{Err}(t)\right)-c_{d+1}\bigg(\tilde{p}(0_{d})A_{M}\tilde{f}(0_{d})+[\tilde{p},\tilde{f}]_{M}(0_{d})\bigg)\right](1+O(1))$
@@
Moreover, the same conclusion as in Theorem 1.1 is true when $x$ is an interior point or a LTCDD point.
It is worth pointing out that in the previous context, the inward tangent sector $I_{x}M$ coincides with the Bouligand tangent cone of $\Omega$ at $0_{d}$, so that Theorem 1.2 matches up with Theorem 1.1. Let us also stress out that, while bounded geometry of the boundary was assumed in [HAL2007], we do not make any assumption of that kind here.
The recent preprint [ART] provides a result that goes in the direction of Theorem 1.2, but with less details on the differential operators involved there. Our approach allows for a neat identification of the latter in several cases, including the following ones.
@@
Theorems 1.1 and 1.2 open the door to a natural Neumann boundary condition on (sub)manifolds with kinks, that would be $\partial_{v_{g}(x)}f(x)=0$ for any $x\in\partial M$. A natural problem to tackle after that is the spectral convergence of the Graph Laplacian with Gaussian kernel generated from finitely many samples $\{X_{1},\ldots,X_{n}\}$ to the suitable Neumann Laplacian as sample size goes to infinity. We refer to [BelkinNiyogi_conv, SingerWu2, TGHS] for related results on manifolds without and with boundary, see also [DT] for a preliminary investigation with the symmetrized AMV kernel.
@@
We obtain Theorem 1.3 in Section 7 through concentration estimates. Such estimates — studied in e.g. [Hein2005, belkin+, AnderssonAvelin] — quantify the error in approximating the Gaussian operator by the graph Laplace operator. Our results generalize this line of work in two key directions :
@@
    We work in the setting of Riemannian manifolds with kinks, whose singularities can be significantly more severe than those of manifolds with boundary considered in [belkin+].
@@ Acknowledgments.
The authors are funded by the Research Foundation – Flanders (FWO) via the Odysseus II programme no. G0DBZ23N. They both thank Dominic Joyce for helpful answers concerning manifolds with corners, Iosif Pinelis who suggested the proof of (7), and Laurent Bessières for a comment on the first version of the paper.
@@ 2. Euclidean models
For a subset $A$ of the Euclidean space $\mathbb{R}^{d}$, we let $\bar{A}$ be the closure of $A$ and $int(A)$ be its interior. We recall that a rigid motion of $\mathbb{R}^{d}$ is a linear map $y\mapsto Ay+b$ where $A\in O(d)$ and $b\in\mathbb{R}^{d}$. The epigraph and strict epigraph of $\gamma:\mathbb{R}^{d-1}\to\mathbb{R}$ are defined, respectively, as
@@
whenever the limit exists. If this is the case, then the second directional derivative of $\gamma$ at $x^{\prime}$ in the direction $v^{\prime}$ is defined as
$\gamma^{\prime\prime}(x^{\prime};v^{\prime}):=\lim_{t\downarrow 0}\frac{\gamma(x^{\prime}+tv^{\prime})-\gamma(x^{\prime})-t\gamma^{\prime}(x^{\prime};v^{\prime})}{t^{2}/2}$
whenever it exists. Lastly, the characteristic function of a subset $A$ of a set $X$ will always be denoted by $1_{A}$.
@@ 2.1. Boundary points
For an open set $\Omega\subset\mathbb{R}^{d}$ with a non-empty boundary $\partial\Omega$, we will call boundary points of $\Omega$ the elements in $\partial\Omega$. Inspired by [Evans, p.626], we define the regularity of boundary points as follows.
@@
    We say that $x$ is a Continuously Directionally Differentiable (CDD for brevity) boundary point if it is a DD point with a function $\gamma$ as above whose directional derivative $\gamma^{\prime}(x^{\prime};v^{\prime})$ exists for any $v^{\prime}\in\mathbb{R}^{d-1}$ and is continuous with respect to $v^{\prime}$. We additionally say that $x$ is TCDD if the second directional derivative $\gamma^{\prime\prime}(x^{\prime};v^{\prime})$ exists for any $v^{\prime}\in\mathbb{R}^{d-1}$ and is continuous with respect to $v^{\prime}$.
As customary, we may write smooth instead of $\mathcal{C}^{\infty}$, and Lipschitz instead of $\mathcal{C}^{0,1}$. We will essentially work with boundary points that are $\mathcal{C}^{0}$, $\mathcal{C}^{1}$, Lipschitz, or CDD. The combination of Lipschitz and CDD (resp. TCDD) shall be abbreviated to LCDD (resp. LTCDD).
@@ 2.2. Diffeomorphisms and inward sectors
For $k$ a positive integer, we shall use the following notion of $\mathcal{C}^{k}$ diffeomorphism between possibly non-open subsets of $\mathbb{R}^{d}$, see e.g. [Melrose, Section 1.5] for details.
@@ 2.3. Boundary cones of open Euclidean subsets
In this section, we provide classical notions from convex analysis, see e.g. [RW] for more details. We recall that $C\subset\mathbb{R}^{d}$ is called a cone if it is invariant under multiplication by a positive number, i.e. if $\lambda v\in C$ for any $v\in C$ and $\lambda>0.$
@@
The Bouligand tangent cone is also known as tangent cone or contingent cone in the literature. It is a closed set. One can easily check that our definition is equivalent to the one given in [RW, Def. 6.1]. Note also that
@@
{subcaptionblock}
.4 Refer to caption {subcaptionblock}.4 Refer to caption
Figure 3.
Figure 4.
@@
{subcaptionblock}
.4 Refer to caption {subcaptionblock}.4 Refer to caption
Figure 6.
Figure 7.
@@ 2.4. Non-fluctuating boundary and local blow-ups
{subcaptionblock}
.4 Refer to caption {subcaptionblock}.4 Refer to caption
Figure 9. Fluctuating $\partial\Omega_{1}$
Figure 10. Non-fluctuating $\partial\Omega_{4}$
@@ 3.1. Topological manifolds with kinks
Recall that a topological space is called paracompact if any open cover has a locally finite open refinement, and that a Hausdorff topological space is paracompact if and only if it admits partitions of unity subordinate to any open cover. For two topological spaces $X,Y$ and a subset $A\subset X$, a map $\varphi:A\to Y$ is a topological embedding if it is an homeomorphism onto its image.
@@ 3.2. Differentiable manifolds with kinks
    We say that $x\in\partial M$ is a $\mathcal{C}^{1}$-boundary point of $M$ if there exists a local border chart $(U,\phi)$ around $x$ such that $\phi(x)$ is a $\mathcal{C}^{1}$ boundary point of $int(\phi(U))$. Otherwise, we say that $x$ is an essential kink.
@@
    We say that an essential kink $x\in\partial M$ is an essential corner of depth $k\in\{2,\ldots,d\}$ if there exists a border chart $(U,\phi)$ centered at $x$ such that, up to a rigid motion, the image $\phi(U)$ locally writes as $\mathbb{R}^{d}_{k}$.
@@
Remark 3.11.
It is worth pointing out that essential corners are Lipschitz border points, as the set $\mathbb{R}^{d}_{k}$ is the epigraph of a Lipschitz function $\gamma:H\to\mathbb{R}$ where $H$ is a (d1)(d-1)-dimensional subspace of $\mathbb{R}^{d}$. Indeed, consider the vector $n=(1,\ldots,1)\in\mathbb{R}^{k}$ and its scaled version $\hat{n}=n/\|n\|\in\mathbb{S}^{k-1}\subset\mathbb{R}^{k}$. Then $P:=\hat{n}^{\perp}$ is a (k1)(k-1)-dimensional subspace of $\mathbb{R}^{k}$, and any $x\in\mathbb{R}^{k}$ writes as $\xi+t\hat{n}$ with $\xi=(\xi_{1},\ldots,\xi_{k-1})\in P$ and $t\in\mathbb{R}$, where the coordinates of $\xi$ are taken with respect to any orthonormal basis of $P$. Therefore
$x=\xi+t\hat{n}\in\mathbb{R}_{k}^{k}\quad\iff\quad t\geq-\sqrt{k}\min_{1\leq i\leq d-1}\xi_{i}=:g_{k,k}(\xi).$
In other words, $\mathbb{R}_{k}^{k}=epi(g_{k,k})$, and $g_{k,k}$ is obviously Lipschitz because it is the pointwise minimum of a family of Lipschitz functions. Setting
$g_{d,k}(\zeta):=g_{k,k}(\pi_{P}(\zeta))\quad\text{ for any $\zeta\in P\oplus\mathbb{R}^{d-k}$},$
where $\pi_{P}$ is the Euclidean orthogonal projection onto $P$, we get that $\mathbb{R}_{k}^{d}=epi(g_{d,k})$, and $g_{d,k}$ is Lipschitz too.
@@
    The pyramid $M:=\{(x,y,z)\in\mathbb{R}^{3}:z\geq\max(|x|,|y|)\}$ is a manifold with kinks that is not a manifold with corner, because the point $0_{3}\in\partial M$ is an essential kink which is not a corner.
@@
{subcaptionblock}
.4 Refer to caption {subcaptionblock}.4 Refer to caption
Figure 13. Pyramid
Figure 14. Cusp
@@ 3.3.1. Tangent space
From now on, we work only with smooth manifolds with kinks, i.e. those manifolds with kinks for which the transition maps are all $\mathcal{C}^{\infty}$. This is to make sure that the tangent space defined below as the space of smooth derivations (i.e. linear forms on the space of smooth functions) is finite dimensional. Indeed, for any $\mathcal{C}^{k}$ manifold with $k<+\infty$, the space of $\mathcal{C}^{k}$ derivations is infinite dimensional : see e.g. [AMR, Theorem 4.2.41].
In order to introduce smooth derivations on manifolds with kinks, we must first define what a smooth function is. This is what we do in the next definition, largely inspired by the context of manifolds with corners (see e.g. [Joyce, Definition 2.2]).
@@
If $x\in\partial M$ and $f\circ\phi^{-1}:\phi(\bar{U})\cap\mathbb{B}_{\rho}^{d}\to\mathbb{R}$ admits Taylor polynomials of arbitrary high order, then the existence of a smooth extension of $f\circ\phi^{-1}$ to $\mathbb{R}^{d}$ is ensured by Whitney’s extension theorem [Whitney, Theorem I], see also [Stein, VI.2.]. Note that in the setting of manifolds with corners, the latter upgrades into the well-posedness of Seeley’s linear extension operator [Seeley], see [Melrose, Section 1.4].
@@
We can now define the differential of a smooth function between smooth manifolds with kinks. There is no change with the case of classical manifolds, see e.g. [Lee, p.55].
@@
When $x$ is an interior point, a $\mathcal{C}^{1}$ boundary point or an essential corner, this is already known, see [Joyce]. Let us then assume that $x$ is an essential kink which is not an essential corner. Consider a local chart $(U,\phi)$ centered at $x$. Like for manifolds without boundary, we obtain from the chain rule applied to the identities $\phi\circ\phi^{-1}=\mathrm{id}_{\phi(U)}$ and $\phi^{-1}\circ\phi=\mathrm{id}_{U}$ that the differential $\mathop{}\!\mathrm{d}_{x}\phi:T_{x}M\to T_{0_{d}}\phi(\bar{U})$ is a linear isomorphism with inverse $\mathop{}\!\mathrm{d}_{0_{d}}\phi^{-1}$. But the space of derivations at the $\mathcal{C}^{0}$ boundary point $0_{d}$ of $\phi(U)$ coincides with $\mathbb{R}^{n}$, since a basis of this space is given by the classical partial differential operators $\partial\cdot/\partial x_{1},\ldots,\,\partial\cdot/\partial x_{n}$ defined on the open set $\phi(U)$ and naturally extended to any open set $V\subset\mathbb{R}^{d}$ containing $\overline{\phi(U)}$. ∎